OpenStudy (anonymous):

Differential equation help. Apparently I am doing something wrong because I did not get the answer in the book can someone spot my mistake? I will write out the problem and what I did below...

OpenStudy (anonymous):

\[x\frac{ dy }{ dx }+y= \frac{ 1 }{ y^2} \] Dividing by x across \[\frac{ dy }{ dx }+\frac{y}{x}= \frac{ 1 }{x y^2} \] Noticing its is a Bernoulli equation and letting N=-2 \[u=y^{3}\] Then, \[y=u^{\frac{1}{3}}\] and \[\frac{ dy }{ dx }= \frac{ 1 }{3}u^{\frac{-2}{3}}\ \] Substituting back in \[\frac{1}{3u^{\frac{2}{3}}}\frac{ du }{ dx }+\frac{u^\frac{1}{3}}{x}= \frac{1}{x\[(u^{\frac{ 1 }{ 3 }}) ^2\]}\ \[(u^{\frac{ 1 }{ 3 }}) \[(u^{\frac{ 1 }{ 3 }}) ^2\]\] (ignore the extra u^1/3 at the end i cant seem to get them to go away lol)

OpenStudy (anonymous):

multiplying through by 3u^(2/3) yeilds: \[\frac{ du }{ dx }+\frac{ 3u}{ x }=\frac{ 3 }{ x }\] finding the integrating factor \[\mu(x)=\[e^{\int\limits_{}^{}\frac{ 1 }{ x }dx}\] = x which leaves \[\frac{ d }{ dx }3ux=3\] solving for u \[u(x)=\frac{ 3x+c }{ 3x } =1+\frac{ c }{ 3x }\]

OpenStudy (anonymous):

apparently the correct answer is \[1+\frac{ c}{ x^{-3} }\]

OpenStudy (lgbasallote):

integrating factor should be \[e^{\int \frac 3x dx}\] right? i don't know if you did some trick afterwards though

OpenStudy (anonymous):

possibly that is my mistake... I'll have to go back and check

OpenStudy (lgbasallote):

yep. if you do that you'll arrive with \[1 + \frac c{x^3}\]

OpenStudy (anonymous):

ah thanks.. always something stupid like that I miss lol.